OpenStudy (anonymous):

Add the following polynomials: (7x2 + x + 13) + (2x2 – 5x + 10)

OpenStudy (anonymous):

you again!!!!!!!! lol

OpenStudy (anonymous):

9x2-4x+23

OpenStudy (anonymous):

7x2 + x + 13 + (2x2 - 5x +10) = 7x2 + x + 13 + 2x2 -5x +10 = (7x2 + 2x2) + (x-5x) + (13+10) = 9x2 - 4x + 23